OpenStudy (anonymous):

For the system: a) Write the coefficient matrix A of the system b) Find det(A) c) Compute |-2A.AT.A-1|

OpenStudy (jdoe0001):

hhmmm for what system?

OpenStudy (anonymous):

X+2Y+Z=3 3X-Y-3Z=-1 2X+3Y+Z=4

OpenStudy (jdoe0001):

the matrix will be all coefficients of the variables arranged in the same order